# Percolation Statistics in the MillenniumTNG Simulations

### v1

@@ 1 Introduction
Figure 1: Percolation statistics for three different Poisson distributions of points, with $256^{3}$ (dashed), $512^{3}$ (dotted) and $1024^{3}$ (solid) particles. For all three cases, we show the mass of the biggest structure ($G_{1}$) in units of the total particle number, as well as the 2nd, 10th, and 100th largest structure as a function of the dimensionless linking length $b$. The percolation transition at $b\simeq 0.865$ is invariant with respect to the point density, as are the absolute particle numbers of the largest structures before the percolation transition occurs, but not when the group sizes are expressed in terms of the fraction of mass they contain, as done here.
@@ 3.1 Random point sets
Note, however, that the full percolation curve is not invariant with respect to particle density (unlike found by Zhang et al., 2018). This is because in the regime prior to percolation the largest group sizes tend to have similar absolute particle numbers for a given characteristic linking length. In the regime the groups do not yet know of the finite periodic domain they live in. However, after percolation, the evolution of the $G_{1}$ curve becomes universal, whereas the shrinking of the other group sizes stays non-universal due to a residual particle density dependence.
@@ 3.3 Neutrinos as hot dark matter
In Figure 6 we show the corresponding results, which interestingly show a substantial difference between the models with $300\,{\rm meV}$ and $100\,{\rm meV}$ neutrinos. The more massive neutrino model shows a substantially earlier percolation transition than the lighter model. The differences are consistent with the expectation that by $z=0$ the more massive neutrinos have begun to cluster more strongly than the lighter ones, which still move much faster. But even in the latter case, the neutrino distribution is clearly very different already from a Poisson distribution, demonstrating that even these light neutrinos have started to develop non-trivial clustering patterns at low redshift, and that this is not drowned in the substantial residual Poisson noise with which the neutrinos are represented in the MTNG N-body simulations. We also include in Fig. 6 measurements of the percolation statistics for the corresponding dark matter particle distributions. Interestingly, even here a very small change in the dark matter clustering due to the different neutrino masses can be detected by the percolation statistics. The result suggests that the clustering of the dark matter in the simulation with the more massive neutrinos ends up being slightly stronger, consistent with the expectation that these neutrinos start to boost structure growth at late times compared with the lighter neutrinos that are still more smoothly distributed.
@@ 4 Percolation of the galaxy distribution
We now turn to considering galaxy mock catalogues produced for our MTNG simulations (Barrera et al., 2023). This is in line with the originally intended application of percolation statistics as a means to quantify the three-dimensional geometry of the galaxy distribution, and to use this for a comparison between the observed cosmic large-scale structure and theoretical predictions. Ideally, such a comparison could be sensitive to features in the cosmic web morphology that are not differentiated by simple two-point clustering statistics.
@@
For definiteness, we consider different number densities of galaxy samples, with values roughly in the ballpark of those of current surveys (the same ones as Angulo et al., 2014). To implement these densities, we rank-order the original full galaxy catalogue for example by stellar mass, and then select the most massive galaxies down to a cut-off value that reproduces the desired space density. This procedure thus mimics typical observational brightness or stellar mass cuts. Besides stellar mass, we also consider total halo mass of the parent halo the galaxy resides in, and the instantaneous star formation rate, as selection criteria. Especially for the SFR cut we expect that the resulting galaxy sample will differ substantially from the one defined through a stellar mass cut, as the latter will also contain many quenched massive galaxies that are disfavoured to be included in the SFR-selected catalogue. In the following, we are especially interested in the question how such expected differences show up in the percolation statistics.
@@
Before we turn to this, it is instructive to first consider the ordinary two-point correlation functions of the different galaxy samples we have defined. They are shown in Figure 7 for the 4 number densities and the 3 selection criteria we consider. The galaxy-two point correlation functions exhibit the familiar shape. Notably, we are able to detect the baryonic acoustic peak at $\sim 120\,{\rm Mpc}$ thanks to the large box size of our simulation. Importantly, the correlation function amplitude of the SFR-selected sample is considerably smaller than that of the stellar mass or halo mass selected samples. Furthermore, the clustering of the different SFR-selected galaxy samples is fairly similar, demonstrating in this case insensitivity to the number density of the considered galaxy sample. In contrast, the clustering amplitude of galaxies becomes systematically higher if rarer samples are selected by stellar or halo mass. This is due to the well-known effect that the galaxy bias increases for more massive halos (Mo & White, 1996). Note that for our mock catalogues, the two-point correlation functions are nearly identical for halo-mass selection or stellar-mass selection. For simplicity, we thus focus in the following on the comparison between halo mass and stellar mass percolation results.
@@
Interestingly, the results of Fig. 11 show that galaxies in redshift space tend to percolate significantly earlier than in real space, and that this effect is more pronounced for our high galaxy number density. Note that the corresponding shifts are similar to – or even considerably higher – than the differences in percolation threshold seen between SFR- and halo-selected samples in real space (Fig. 10). Again, this is a fairly sobering result for the discriminative power of percolation statistics as a practical tool in cosmology. Note also that our results do not confirm earlier findings by Zhang et al. (2018) who reported an insensitivity of percolation statistics to redshift space distortions.
@@ 5 Discussion
Furthermore, it turns out that redshift space distortions have a quite substantial effect on the percolation statistics, lowering the percolation threshold, probably simply due to the finger-of-god effect. It is thus unfortunately not true that redshift space distortions can be neglected for percolation statistics, at least not at the galaxy sample densities that we have considered here.
Our results thus support the cautionary warnings of Dekel & West (1985) in early work on percolation statistics. A meaningful application of percolation statistics as a cosmological test seems only be possible if the number densities of observational and theoretical samples are matched. In addition, one needs to consistently work in redshift or real space. Then, however, percolation can still provide a useful test for checking whether the large-scale geometry of the cosmic web is consistent or not. In this function it can complement other topological and morphological measures that are sensitive to higher-order correlation functions.
@@ 6 Conclusions
  • Redshift space distortions tend to lower the percolation threshold at all galaxy densities. The differences between real space and redshift space tend to be sizeable, and typically larger than variations induced by different galaxy selection criteria.
@@ Acknowledgements
For carrying out the MillenniumTNG simulations, the authors gratefully acknowledge the Gauss Centre for Supercomputing (GCS) for providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (LRZ) in Garching, Germany, under project pn34mo. The work also used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility, with equipment funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. ER is supported by NKFIH-OTKA grant K-142534 from the National Research, Development and Innovation Office, Hungary. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. CH-A acknowledges support from the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@
The simulations of the MillenniumTNG project are foreseen to be made fully publicly available in 2024 at the following address: https://www.mtng-project.org. The data underlying this article is available upon reasonable request to the corresponding author.

### v2

@@ 1 Introduction
Figure 1: Percolation statistics for three different Poisson distributions of points, with $256^{3}$ (dashed), $512^{3}$ (dotted) and $1024^{3}$ (solid) particles. For all three cases, we show the mass of the biggest structure ($G_{1}$) in units of the total particle number, as well as the 2nd, 10th, and 100th largest structure as a function of the dimensionless linking length $b$. The percolation transition at $b\simeq 0.865$ is invariant with respect to the point density, as are the absolute particle numbers of the largest structures before the percolation transition occurs. However, the percolation curve depends on the particle number when the group sizes are expressed in terms of the fraction of mass they contain, as done here.
@@ 3.1 Random point sets
Note, however, that the full percolation curve is not invariant with respect to particle number (unlike found by Zhang et al., 2018). This is because in the regime prior to percolation the largest group sizes tend to have similar absolute particle numbers for a given characteristic linking length. In the regime the groups do not yet know of the finite periodic domain they live in. However, after percolation, the evolution of the $G_{1}$ curve becomes universal, whereas the shrinking of the other group sizes stays non-universal due to a residual particle density dependence.
@@ 3.3 Neutrinos as hot dark matter
In Figure 6 we show the corresponding results, which interestingly show a substantial difference between the models with $300\,{\rm meV}$ and $100\,{\rm meV}$ neutrinos. The more massive neutrino model shows a substantially earlier percolation transition than the lighter model. The differences are consistent with the expectation that by $z=0$ the more massive neutrinos have begun to cluster more strongly than the lighter ones, which still move much faster. But even in the latter case, the neutrino distribution is clearly very different already from a Poisson distribution, demonstrating that even these light neutrinos have started to develop non-trivial clustering patterns at low redshift, and that this is not drowned in the substantial residual Poisson noise with which the neutrinos are represented in the MTNG N-body simulations. We also include in Fig. 6 measurements of the percolation statistics for the corresponding dark matter particle distributions. We expect that the clustering of the dark matter in the simulation with the more massive neutrinos ends up being slightly stronger, consistent with the expectation that these neutrinos start to boost structure growth at late times compared with the lighter neutrinos that are still more smoothly distributed. Our percolation statistics results seem to indicate a very small systematic shift consistent with this expectation, but the effect is so small that it is not securely detected above the measurement noise.
@@ 4 Percolation of the galaxy distribution
We now turn to considering galaxy mock catalogues produced for our MTNG simulations (Barrera et al., 2023). This is in line with the originally intended application of percolation statistics as a means to quantify the three-dimensional geometry of the galaxy distribution, and to use this for a comparison between the observed cosmic large-scale structure and theoretical predictions. Ideally, such a comparison could be sensitive to features in the cosmic web morphology that are not differentiated by simple two-point clustering statistics. Note that unlike for the point sets analysed so far, the galaxy distribution can be affected by small-scale exclusion effects as galaxies cannot freely overlap due to their finite size. For the galaxy densities we will consider, their mean distance is far larger than their average size, so that the corresponding effects are weak.
@@
For definiteness, we consider different number densities of galaxy samples, with values roughly in the ballpark of those of current surveys (the same ones as Angulo et al., 2014). To implement these densities, we rank-order the original full galaxy catalogue for example by stellar mass, and then select the most massive galaxies down to a cut-off value that reproduces the desired space density. This procedure thus mimicks typical observational brightness or stellar mass cuts. Besides stellar mass, we also consider total halo mass of the parent halo the galaxy resides in, and the instantaneous star formation rate, as selection criteria. Especially for the SFR cut we expect that the resulting galaxy sample will differ substantially from the one defined through a stellar mass cut, as the latter will also contain many quenched massive galaxies that are disfavoured to be included in the SFR-selected catalogue. In the following, we are especially interested in the question how such expected differences show up in the percolation statistics.
@@
Before we turn to this, it is instructive to first consider the ordinary two-point correlation functions of the different galaxy samples we have defined. They are shown in Figure 7 for the 4 number densities and the 3 selection criteria we consider. The galaxy-two point correlation functions exhibit the familiar shape. Notably, we are able to detect the baryonic acoustic peak at $\sim 120\,{\rm Mpc}$ thanks to the large box size of our simulation. Importantly, the correlation function amplitude of the SFR-selected sample is considerably smaller than that of the stellar mass or halo mass selected samples. Furthermore, the clustering of the different SFR-selected galaxy samples is fairly similar, demonstrating in this case insensitivity to the number density of the considered galaxy sample. In contrast, the clustering amplitude of galaxies becomes systematically higher if rarer samples are selected by stellar or halo mass. This is due to the well-known effect that the galaxy bias increases for more massive halos (Mo & White, 1996). Note that for our mock catalogues, the two-point correlation functions are nearly identical for halo-mass selection or stellar-mass selection. For simplicity, we thus focus in the following on the comparison between halo mass selected and star formation rate selected percolation results.
@@
Interestingly, the results of Fig. 11 show that galaxies in redshift space tend to percolate significantly earlier than in real space, and that this effect is more pronounced for our high galaxy number density. Note that the corresponding shifts are similar to – or even considerably higher – than the differences in percolation threshold seen between SFR- and halo-selected samples in real space (Fig. 10). Again, this is a fairly sobering result for the discriminative power of percolation statistics as a practical tool in cosmology. Note also that our results do not confirm earlier findings by Zhang et al. (2018) who reported an insensitivity of percolation statistics to redshift space distortions. However, their galaxy sampling density is comparatively low, a regime where we also find only a small impact of redshift space distortions.
@@ 5 Discussion
Furthermore, it turns out that redshift space distortions have a quite substantial effect on the percolation statistics, lowering the percolation threshold, probably simply due to the finger-of-god effect. It is thus unfortunately not true that redshift space distortions can be neglected for percolation statistics, at least not at the comparatively high galaxy sample densities that we have considered here.
Figure 12: Percolation statistics for halo samples of constant space density equal to $n=1.0\times 10^{-3}\,{\rm Mpc}^{-3}$, selected by virial mass at $z=0$, in four different simulations with the same cosmological model. We compare two box sizes, equal to $L=3000\,{\rm Mpc}$ and $L=632\,{\rm Mpc}$, each carried out with two paired realizations A and B, as labelled. The thick lines give the average of the two runs for the corresponding box size, while the thin lines show the individual simulations. Clearly, the large box produces results that are hardly affected by cosmic variance at all, and even the small box shows only small differences between the realizations A and B. Note, however, that prior to the percolation statistics, the results for $G_{1}(b)$ do depend not only on the sampling density but also on the box size. However, if one defines the percolation transition through, for example $G_{1}(b)=0.25$ (dotted line), nearly invariant results for the percolation threshold are obtained.
Another interesting question is whether cosmic variance is a significant issue for our results, given that we have thus far analyzed individual simulation realizations of specific cosmological models. Note, however, that our volumes have been comparatively large, especially for our galaxy catalogues, which were constructed for a simulation box 3000 Mpc on side. This is usually considered to be large enough to give results that are fully representative for the underlying cosmological model. A case in point is the good reproduction of the baryonic acoustic oscillation feature in the two-point correlation function seen in Figure 7. But to firm this up more, we consider in Figure 12 the percolation statistics for both of the two simulations we have carried out in the MTNG project for the $3000\,{\rm Mpc}$ boxsize. These realizations differ in the sign of their initial linear displacement field, so that their cosmic web ends up being complementary to each other, in the sense that clusters tend to become voids and vice versa in the evolved simulations. In addition, we also include results for two simulations in a much smaller box of $632\,{\rm Mpc}$ on a side, at the same mass resolution and using the same cosmological model (i.e. the particle number and volume of these simulations are more than 100 times smaller than for the big boxes). Again, we consider two realizations of this smaller box. For definiteness, we consider the most massive halos selected by virial mass with space density $n=1.0\times 10^{-3}\,{\rm Mpc}^{-3}$ in all the cases.
As Figure 12 shows, the results of the two simulation realizations are very similar in both cases, with the smaller boxes showing slightly larger cosmic variance scatter whereas this is essentially negligible for the big box. Note, however, that prior to the percolation transition, the percolation curve does depend on the box size (or equivalently on the total number of halos) even though the sample density is the same – this is the same effect that we already noticed for random point distributions. However, the percolation transition occurs at the same place for both box sizes, and if one defines this implicitly through, for example, $G_{1}(b)=0.25$ one obtains results for the percolation threshold that are independent of box size to good accuracy.
Our results thus support the cautionary warnings of Dekel & West (1985) in early work on percolation statistics. A meaningful application of percolation statistics as a cosmological test seems only be possible if the number densities and volumes of observational and theoretical samples are matched. In addition, one needs to consistently work in redshift or real space. Then, however, percolation can still provide a useful test for checking whether the large-scale geometry of the cosmic web is consistent or not. In this function it can complement other topological and morphological measures that are sensitive to higher-order correlation functions.
@@ 6 Conclusions
  • Redshift space distortions tend to lower the percolation threshold at all galaxy densities. The differences between real space and redshift space tend to be sizable, and typically larger than variations induced by different galaxy selection criteria.
@@ Acknowledgements
We thank the anonymous referee for constructive and insightful comments that helped to improve the paper. For carrying out the MillenniumTNG simulations, the authors gratefully acknowledge the Gauss Centre for Supercomputing (GCS) for providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (LRZ) in Garching, Germany, under project pn34mo. The work also used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility, with equipment funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. ER is supported by NKFIH-OTKA grant K-142534 from the National Research, Development and Innovation Office, Hungary. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. CH-A acknowledges support from the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@
The simulations of the MillenniumTNG project are foreseen to be made fully publicly available in 2024 at the following address: https://www.mtng-project.org.